Researcher is using data for a sample of 10 observations to estimate the relation 

between consumption expenditure and income. Preliminary analysis of the sample data 

produces the following data. 

 ∑xy = 700 , 1000 2 ∑x = , ∑X = 100

 ∑Y = 200 

Where i x Xi X

__

= − and 

__

y = Yi − Y

a. Use the above information to compute OLS estimates of the intercept and slope 

coefficients and interpret the result 

b. Calculate the variance of the slope parameter 

c. Compute the value R

2

 (coefficient of determination) and interpret the result 

d. Compute 95% confidence interval for the slope parameter 

e. Test the significance of the slope parameter at 5% level of confidence using t-test


1
Expert's answer
2022-02-01T08:54:25-0500

(a) Define Y^ = b0 + b1x

where b0 is the y intercept and b1 is the slope or regression coefficient of the line.

b1 =  (∑xy - (∑X ∑Y)/n ) /( ∑X2 - ( ∑X )2/n )

which yields

= (700 - (100 * 200)/10) / ( 1000 - (100)2/10)

= (-1300)/ ( 0 ) which is undefined

since b1 is undefined,it makes it impossible to handle part b, c, d and e respectively.


b) variance of the slope = MSE/ ∑( XI - X- )2


c) coeeficient of determination = ( coefficient of correlation)2


d) t = b1/sb1 , where b1 is the slope and sb1 is the "square\,root" of the variance of the slope


e) if tcalculated > ttabulated ,then we conclude that the slope is significant.